OpenStudy (anonymous):

There are 40 elephants. 6 are blue, 8 black, 10 red, 6 yellow and 10 green. The 40 elephants are raced. Given that there are exactly 4 red in the first 10 place what is the probability that there are 4 blue also

OpenStudy (mathmate):

Use conditional probability P(A|B)=P(A intersect B)/P(B)

OpenStudy (anonymous):

we know that there are 4 red elephants accounted for, that leaves \(40-4=36\) elephants in our sample space now we want the probability that 4 out of the 6 in first place are blue

OpenStudy (anonymous):

lost the connection

OpenStudy (anonymous):

number of ways to pick 6 out of 36 is \[\dbinom{36}{6}\] you want 4 blue and 2 not blue and not red there are 6 blue total, the number of ways to pick 4 out of 6 is \[\dbinom{6}{4}=\frac{6\times 5}{2}=15\]

OpenStudy (anonymous):

the number of ways to pick 2 out the remaining not blue and not red elephants is \[\dbinom{24}{2}=\frac{24\times 23}{2}=12\times 23=276\]

OpenStudy (anonymous):

check my arithmetic here because i sort of lost my train of though during the blackout in any case i believe you need to compute \[\frac{\dbinom{6}{4}\times \dbinom{24}{2}}{\dbinom{36}{6}}\]

OpenStudy (anonymous):

notice that in a conditional probability question like this one, you use what you know to narrow down the options

OpenStudy (anonymous):

i tried it using conditional prob.. P(4B|4R) = (6C4x10C4x24C2 / 40C10) / (10C4x30C6/40C10) which cancels to (6C4x24C2) / 36C6.. is it wrong

OpenStudy (anonymous):

(6C4x25C2) / 30C6 i mean

OpenStudy (anonymous):

Given that there are exactly 4 red in the first 10 place we know that there are 4 red elephants accounted for, that leaves 40−4=36 elephants in our sample space doesnt exactly 4 red means that red cant be used anymore so 40-10=30 ?

OpenStudy (anonymous):

oh maybe i was wrong

OpenStudy (anonymous):

im not sure..

OpenStudy (anonymous):

when you say "it is wrong" do you know the answer?

OpenStudy (anonymous):

nope sorry this wass on the test.. left it blank though

OpenStudy (anonymous):

ok let me try again

OpenStudy (anonymous):

probability that 4 are red AND 4 are blue is \[\frac{\dbinom{10}{4}\dbinom{6}{4}\dbinom{24}{2}}{\dbinom{40}{10}}\]

OpenStudy (anonymous):

probability that 4 are red is \[\frac{\dbinom{10}{4}\dbinom{30}{6}}{\dbinom{40}{4}}\]

OpenStudy (anonymous):

yes i was wrong divide and get \[\frac{\dbinom{6}{4}\dbinom{24}{2}}{\dbinom{30}{6}}\]

OpenStudy (anonymous):

you are right, the ten red ones are out of the picture that is what i get for trying to think instead of using the formula

OpenStudy (anonymous):

so what you wrote above (6C4x25C2) / 30C6 is right, except i think you meant 24 and not 25

OpenStudy (anonymous):

yup.. next part.. 3 of the yellow elephants and tied together whilst 4 of the blue are tied toegher, what is the probability that these 7 elephants are in top 10?

OpenStudy (anonymous):

hi asdhsadas